A bag contains 4 blue chips and 6 red chips. Chi and Dolly play the following game. Chi selects one chip at random from the bag. If Chi selects a blue chip, Dolly gives Chi $6. If Chi selects a red chip Chi gives Dolly $5.

Determine Chi's expectation

Determine Dolly's expectation
Answer #1

Probability of blue chip = 4/10 = 0.4

Probability of red chip = 6/10 = 0.6

a) Chi's expectation = 0.4*6 + 0.6 *(-5) = $ -0.6

b) Dolly's expectation = 0.4*(-6) + 0.6 *(5) = $ 0.6